### v2

# Accuracy and Political Bias of News Source Credibility Ratings by Large Language Models

Abstract
Search engines increasingly leverage large language models (LLMs) to generate direct answers, and AI chatbots now access the Internet for fresh data. As information curators for billions of users, LLMs must assess the accuracy and reliability of different sources. This paper audits eight widely used LLMs from three major providers—OpenAI, Google, and Meta—to evaluate their ability to discern credible and high-quality information sources from low-credibility ones. We find that while LLMs can rate most tested news outlets, larger models more frequently refuse to provide ratings due to insufficient information, whereas smaller models are more prone to hallucination in their ratings. For sources where ratings are provided, LLMs exhibit a high level of agreement among themselves (average Spearman’s $\rho=0.81$), but their ratings align only moderately with human expert evaluations (average $\rho=0.59$). Analyzing news sources with different political leanings in the US, we observe a liberal bias in credibility ratings yielded by all LLMs in default configurations. Additionally, assigning partisan identities to LLMs consistently results in strong politically congruent bias in the ratings. These findings have important implications for the use of LLMs in curating news and political information.
Code & datasetshttps://github.com/osome-iu/llm˙domain˙rating
Introduction
Large language models (LLMs) are being integrated into our information ecosystems, transforming how people seek and consume information. One significant trend is the emergence of AI-augmented search, where an LLM layer is added into traditional search engines to provide direct answers based on search results (Xiong et al. 2024). Major platforms like Google111blog.google/products/search/generative-ai-google-search-may-2024 and Microsoft222blogs.microsoft.com/blog/2023/02/07/reinventing-search-with-a-new-ai-powered-microsoft-bing-and-edge-your-copilot-for-the-web have implemented this feature, and newer products like Perplexity AI and You.com have gained substantial user bases and investments.333techcrunch.com/2024/01/04/ai-powered-search-engine-perplexity-ai-now-valued-at-520m-raises-70m Additionally, AI chatbots are now often connected to the Internet, allowing them to fetch up-to-date information not included in their training data and ground their responses in real-time (Vu et al. 2023). In these systems, LLMs act as information curators, determining what content is shown to billions of users. Recent studies suggest that such integration lowers the information access barrier (Wu et al. 2020) and enables users to perform complex tasks more quickly (Spatharioti et al. 2023; Suri et al. 2024), indicating potential for mainstream adoption.
Despite such advantages, the additional LLM components could introduce new problems (Memon and West 2024) because language models face critical technical challenges, such as hallucinations (Ji et al. 2023), and exhibit various biases (Gallegos et al. 2024). Recent audits of popular AI search engines reveal that their results often contain unsupported claims (Liu, Zhang, and Liang 2023) and biases depending on the queries (Li and Sinnamon 2024). Furthermore, experiments conducted by Sharma, Liao, and Xiao (2024) demonstrate that users may engage with more biased information when interacting with AI search engines, and LLMs with predefined opinions can exacerbate this bias. Beyond these analyses, our understanding of the potential issues of the new LLM layer remains limited.
In this paper we wish to quantify the accuracy with which LLMs can determine the credibility of information sources, a crucial qualification for information curators. To this end, we present experiments that audit eight widely used LLMs from three major providers: OpenAI, Meta, and Google. These models are instructed to provide ratings for over 7,000 websites representing major news sources on the Internet. The accuracy of these ratings is measured by their alignment with human expert evaluations.
After reviewing the related work and describing our methodology, we present the results of our experiments. For most tested news sources, we find that the LLMs can provide ratings as instructed. However, larger models more frequently refuse to rate less popular sources due to insufficient knowledge, while smaller models are more prone to hallucinating responses. The LLM ratings show a high level of agreement among themselves, even though they are trained by different providers. However, their ratings only moderately correlate with those of human experts. Focusing on news sources with clear political leanings in the US, we find that assigning partisan identities to LLMs consistently biases their ratings in favor of sources with congruent leanings. Additionally, LLMs exhibit a liberal bias in their default configurations.
Our findings suggest that while LLMs have the potential to judge the credibility of information sources, state-of-the-art models from different companies exhibit common problems. Their lack of knowledge about unpopular information sources presents challenges in dealing with data voids (Boyd and Golebiewski 2018; Aslett et al. 2024). Furthermore, inaccurate LLM ratings due to issues like hallucinations and biases could inadvertently amplify misinformation while suppressing high-credibility sources. When handling political information, the inherent partisan biases in LLMs could exacerbate echo chambers and polarization (Cinelli et al. 2021; Conover et al. 2021). Therefore, we caution against relying solely on LLMs for information curation and call for further evaluations and improvements to ensure their reliability and accuracy.
Related Work
Our study follows a long line of scholarship aimed at understanding how online platforms affect user information diets (Rahwan et al. 2019) and their contributions to issues such as the proliferation of misinformation (Lazer et al. 2018), echo chambers (Cinelli et al. 2021), and polarization (Conover et al. 2021). As prominent platforms for information-seeking, search engines have been scrutinized by many researchers evaluating if their ranking algorithms favor certain news sources (Robertson et al. 2018; Trielli and Diakopoulos 2019). Social media platforms are another focus of attention. Previous studies address topics such as political biases and the amplification of malicious content in various platforms, including Twitter/X (Chen et al. 2021), Facebook (Bakshy, Messing, and Adamic 2015), and YouTube (Hosseinmardi et al. 2021).
The emergence of LLMs such as ChatGPT has initiated a new line of research auditing these models to ensure they are deployed in ways that are ethical, safe, and accountable (Mökander et al. 2023). Recent studies show that LLMs can perpetuate discrimination due to their inherent biases (Bai et al. 2024; Salinas et al. 2023) and stereotypes (Abid, Farooqi, and Zou 2021; Cheng, Durmus, and Jurafsky 2023). In addition, LLMs have often been found to exhibit a liberal bias in the political context (Hartmann, Schwenzow, and Witte 2023; Santurkar et al. 2023).
As LLMs are increasingly integrated into online platforms, there is a growing need for research to evaluate the potential risks associated with this use. Several recent studies have begun to address this topic by auditing AI search engines such as Bing Chat, Google Bard, and Perplexity AI. By querying these tools with various search phrases, Liu, Zhang, and Liang (2023) show that the responses often contain unsupported claims (Liu, Zhang, and Liang 2023); Li and Sinnamon (2024) identify sentiment and geographic biases; and Urman and Makhortykh (2023) reveal significant disparities across different services when dealing with political information. Through experiments in which participants interact with AI search engines in a lab environment, Sharma, Liao, and Xiao (2024) find that users tend to engage with more biased information, and opinionated LLMs can exacerbate this bias. Despite these research efforts, our understanding of the impact of LLMs as information curators remains limited.
Methods
Model Selection
We audit eight state-of-the-art models from three major providers. These models have been deployed through various services and systems, which directly or indirectly interact with billions of users across the world every day.
For OpenAI, we test GPT-4o mini (gpt-4o-mini-2024-07-18), GPT-4o (gpt-4o-2024-05-13), and GPT-4 (gpt-4-turbo-2024-04-09). Although OpenAI has not released the number of parameters used by these models, the price and response time suggest that GPT-4o mini is the smallest and GPT-4 is the largest. These models are available to users through ChatGPT and serve as the backend for AI search engines like Bing. In our experiment, we query OpenAI’s API endpoints directly.
For Meta, we test Llama 3.1 models with 8B, 70B, and 405B parameters (Llama Team, AI at Meta 2024). Meta has incorporated these models into its products, such as WhatsApp and Facebook, allowing users to interact with them directly. Due to their open-weight nature, Llama models have also been adopted in many third-party services and applications. Here, we query the Llama 3.1 models hosted by together.ai through its API endpoints.
For Google, we test Gemini 1.5 Flash (gemini-1.5-flash-001) and Pro (gemini-1.0-pro-001), which are available to users through the Gemini App. Google also leverages these models to provide AI summaries in its search results. We query Google API endpoints directly and disable all safety filters.
Prompts
For all models, we set the temperature parameter to zero and use the same prompts. The system prompt reads as follows:
System prompt You are an assistant to determine the credibility of websites.
Then we provide the model with the task instructions:
Task instruction Rate the website’s credibility: {domain}, on a scale between 0 and 1, where 0 means very low credibility and 1 means very high credibility. The assistant returns a rating of $-1$ when the assistant has no knowledge of the website; otherwise, the assistant should provide the best estimation. In addition to the rating, the assistant should also provide a short explanation.
In each query, {domain} is replaced with the news website of interest. Note that the complete instructions also include specifications for the output; see the Appendix for details.
When evaluating the impact of assigning political identities to LLMs, we append the following instruction to the system prompt:
Political identity assignment You identify as {identity} on the US political spectrum.
In the experiments, {identity} is replaced with the one of “a Democrat,” “an Independent,” or “a Republican” to induce personas with different political views.
News Outlet Credibility Ratings
In this study, we adopt ratings of news source credibility compiled by Lin et al. (2023). The authors analyze the news ratings from six sources, including NewsGuard,444newsguardtech.com/ratings/rating-process-criteria adfontesmedia.com, Iffy index of unreliable sources,555iffy.news MBFC,666mediabiasfactcheck.com/methodology and two lists compiled by professional fact-checkers and researchers (Pennycook and Rand 2019; Lasser et al. 2022). The comparison of these ratings reveals a high level of consistency. Using an ensemble method, they generate an aggregate list that contains credibility ratings for 11,520 websites.
Website Popularity Ranking
We adopt the Tranco list to measure website popularity (Pochat et al. 2018). This list combines the website ranking information from multiple sources, including Alexa777alexa.com/topsites and Cisco Umbrella.888umbrella-static.s3-us-west-1.amazonaws.com/index.html The Tranco list is updated on a routine basis, and for this study we use the snapshot from September 2022. This snapshot contains the ranks of the top 5.5 million websites worldwide. Websites not on the list typically have limited viewership or have been deactivated.
Merging the Tranco list with the list from Lin et al. (2023), we end up with 7,523 websites with credibility ratings and popularity ranks. In the following, we refer to these as the “human expert ratings” and use them as a reference to measure the accuracy of LLM ratings.
Source Political Bias Score
To quantify partisan bias of the news sources in the US political context, we leverage the scores generated by Robertson et al. (2018). The authors first construct a panel of over half a million Twitter accounts matched with US voter registration records and then measure the partisan bias of the audience of each source. The results are scores between $-1$ and $+1$, where a score of $-1$ means that a source is shared exclusively by Democrats, and a score of $+1$ means that a source is shared exclusively by Republicans. The authors show that their scores correlate highly with other ratings, such as the ones produced by Bakshy, Messing, and Adamic (2015), Budak, Goel, and Rao (2016), and allsides.com. Merging the list from Robertson et al. (2018) with the human expert ratings of credibility, we obtain 2,683 sources with political leanings to analyze the political biases of LLMs.
Results
LLM Response Analysis
As described in the Methods section, we first run the 7,523 news sources through all eight LLMs using the same prompt in the default configuration (no assigned political identities). In most cases, the LLMs are able to respond with the required content in the specified format. When errors occur, we rerun the queries until we obtain the desired outputs.
Take reuters.com as an example. GPT-4o provides the following response:
GPT-4o’s assessment of reuters.com Rating: 1.0 Explanation: Reuters is a well-established international news organization known for its accurate and unbiased reporting. It is widely respected in the journalism community and has a long history of providing reliable news.
All other models give Reuters credibility scores over 0.9 with similar explanations (complete responses available in the Appendix). These responses indicate that LLMs recognize news outlets from their websites, possess information about them, and can provide credibility ratings.
Figure 1: Relationship between source popularity and the responses of LLMs. The left axes and the lines show the percentages of sources for which each LLM provides ratings. The dashed lines indicate the overall percentages, whereas the solid lines illustrate the results for different source ranking deciles. The right axes and the dots represent the Spearman correlation coefficients between LLM ratings and human expert ratings in different source ranking deciles. Sources in larger ranking deciles are less popular.
When the LLMs lack sufficient information about a given source, they respond with a rating of $-1$, as instructed. In Figure 1, we illustrate the percentage of sources for which each LLM provides ratings (dashed lines). We find that within each family, larger models are more likely to indicate insufficient information about the given sources and to refuse to rate them. We hypothesize that LLMs tend to lack knowledge about unpopular news sources. To confirm this, we split the sources into 10 deciles based on the popularity ranking and calculate the percentage of sources with LLM ratings within each decile. The solid lines in Figure 1 support our hypothesis for all models.
Figure 2: Percentage of responses based on hallucinations among 200 manually annotated cases for each LLM.
Figure 1 also shows that Llama models can provide ratings for more sources compared to GPT and Gemini models. However, LLMs are known to suffer from hallucinations, where they generate baseless responses to user requests (Ji et al. 2023). To gain deeper insights into the ratings generated by these LLMs, we randomly select 200 sources and manually analyze the responses from all eight models. Through this process, we identify two common types of hallucinations. First, the LLMs can mistakenly associate a website with the wrong organization. For example, five out of eight LLMs identify aldf.com, the official website of the “American Lyme Disease Foundation,” as belonging to the “Animal Legal Defense Fund” (aldf.org). The complete responses of the LLMs can be found in the Appendix. Second, the LLMs can fail to identify the websites yet still generate ratings for them.
In Figure 2, we illustrate the percentage of unambiguous hallucinations among the annotated cases for each LLM. The results show that smaller models tend to hallucinate more frequently within each family. Across different providers, Llama and Gemini models exhibit higher rates of hallucination. It is important to note that even when the models are not hallucinating, they may still provide inaccurate ratings for the sources due to other issues. Next, we assess the accuracy of the LLM ratings by comparing them with those given by human experts.
Rating Accuracy
Figure 3: Heatmap of source credibility rating correlation (Spearman) among different LLMs and human experts. Results in the upper right triangle of the heatmap are based on 3,606 (48.0%) sources rated by all LLMs. Results in the lower left triangle are based on the sources rated by both raters in comparison.
Let us measure how well the LLM ratings correlate with each other and align with those from human experts. For each pair of raters (LLMs or human experts), we calculate the Spearman rank correlation coefficient $\rho$ between their ratings. Given that some LLMs only provide ratings for a subset of the sources, we use two different approaches to measure the correlations. First, we include only the 3,606 (48.0%) sources that are rated by all LLMs. Second, for each pair of raters, we focus on their intersection. The results from both approaches are shown in Figure 3 and demonstrate similar patterns. Therefore, our discussion below will focus on the results from the first approach.
All correlation coefficients in Figure 3 are positive and statistically significant with $p<0.001$. We find a high agreement level among LLMs, with an average $\rho=0.81$, despite their different providers. Conversely, they only moderately correlate with human experts, with an average $\rho=0.59$ and small variation across models.
We also test the impact of website popularity on the accuracy of LLM ratings by splitting the sources into 10 popularity ranking deciles and calculating rating correlations between each LLM and human experts. The results, shown as dots in Figure 1, do not indicate a clear association between LLMs accuracy and source popularity.
Political Biases
To probe political biases in LLM ratings, we focus on the 2,683 sources with partisan leanings in the US context. As mentioned in the Methods section, we query the LLMs with the same prompts except for assigning them different political identities. Specifically, for each source, we instruct the models to rate its credibility from the viewpoints of Democrats, Independents, and Republicans, respectively.
Figure 4: Heatmaps of Spearman correlation coefficients among the ratings generated by LLMs with different political identities. The highest correlation coefficients between the default LLM configuration and different political identities are highlighted by squares with solid edges. The highest correlation coefficients between human experts and different political identities are highlighted by squares with dashed edges.
In Figure 4, we show the correlations between the ratings generated by LLMs with different political identities and those from human experts. Ratings from LLMs in the default configuration (no political identities) are closest to those from the Independent identities, followed by Democratic identities. The correlations with the LLMs identified as Republicans are the lowest. When compared with human experts, the LLMs identified as Independents show higher correlations than all other cases, including the default ones. These patterns are consistent across all models.
To quantify the political biases of LLMs with different political identities, we calculate the LLM rating bias score for each source, defined as the difference between the LLM rating and the human expert rating. This metric accounts for the fact that left-leaning sources in our dataset tend to have higher human expert ratings. A positive/negative bias score means the LLM considers the source more/less credible than expected.
Figure 5: Distributions of LLM rating bias scores of GPT-4o mini with different political identities. The blue and red violins represent the results for left- and right-leaning sources, respectively. Significance of t-tests is indicated by ***: $p<0.001$, *:$p<0.05$, NS: not significant.
Based on the scores provided by Robertson et al. (2018), we classify the 2,683 sources as left- or right-leaning using zero as the threshold and compare their LLM rating bias scores. Note that Chen et al. (2021) suggest using 0.058 as the threshold, which leads to qualitatively similar results in our experiments according to robustness checks.
In Figure 5, we show the distributions of LLM rating bias scores for GPT-4o mini with different political identities. We find that the default configuration and the Democratic identity exhibit liberal bias and are more likely to assign higher-than-expected credibility scores to left-leaning sources. The Republican identity, on the other hand, favors right-leaning sources. The Independent identity shows no significant differences in LLM rating bias scores between left- and right-leaning sources.
Figure 6: Political biases of different LLM-identity configurations. The political biases are measured by t-statistics between the distributions of LLM rating bias scores on left- and right-leaning sources. A negative/positive t-statistic means the LLM-identity configuration favors left-/right-leaning news outlets. The solid symbols indicate statistically significant differences ($p<0.05$).
We replicate the analysis for all LLMs and show the results in Figure 6. Due to space constraints, we only report the t-statistics from the tests comparing the distributions of LLM rating bias scores for left- and right-leaning news sources. For all eight models, assigning a Democratic identity leads to a liberal bias, whereas assigning a Republican identity leads to a conservative bias. The Independent identity shows no significant political biases for half of the models but a liberal bias for the other half, although less pronounced than the Democratic identity. Under the default configuration, where no explicit political identities are assigned to the LLMs, all models except for GPT-4 show a statistically significant liberal bias.
Political Bias and Rating Accuracy
Figure 7: Political bias vs. rating accuracy for all LLM-identity configurations. We use the t-statistics from the tests comparing the distributions of LLM rating bias scores on left- and right-leaning sources to quantify the political bias, and the correlation with human experts for accuracy. LLM-identity configurations with left or right biases are separated and the lines represent linear regressions on the two groups.
The results in Figures 4 and 6 suggest a negative correlation between the political biases and the accuracy of LLM-identity configurations. To confirm this, we present a scatter plot of all model-identity configurations in Figure 7. Regardless of the direction, we find that a stronger political bias correlates with lower rating accuracy. This finding suggests that the misalignment between LLMs and human experts is partially due to the political biases embedded in LLMs, and that reducing these biases could enhance rating accuracy.
Pennycook and Rand (2019) ran an experiment in which participants produced news source ratings that correlated with their different ideologies. However, the combined ratings from participants on both sides correlate strongly with expert ratings. Given that LLMs exhibit consistent biases congruent with assigned identities, we test whether combining ratings from LLMs with different viewpoints could reduce overall biases and improve accuracy.
We consider two aggregate ratings for each model: (1) D+R: the average ratings of the Democratic and Republican identities; and (2) D+I+R: the average ratings of the Democratic, Independent, and Republican identities. Although different weights can be assigned to these political identities, we choose equal weights to reflect the composition of American voters with varying partisanship.999pewresearch.org/politics/2024/04/09/the-partisanship-and-ideology-of-american-voters
Figure 8: Accuracy of different LLM ratings. We show the Spearman correlation coefficients between different LLM ratings and those from human experts. For each model, we report the results for the default configuration, three political identities (Democrat, Independent, and Republican), as well as two aggregate ratings, D+R (average of Democratic and Republican identities) and D+I+R (average of all three political identities). The top three configurations for each model are annotated.
In Figure 8, we compare the accuracy of D+R and D+I+R with the results for different political identities. In most cases, D+R and D+I+R ratings are more accurate than either Democratic or Republican ratings, suggesting that blending different perspectives leads to better judgment of source credibility. Across all configurations, the Independent identity is the winner for four models, while D+I+R leads in the other four cases. However, the differences between the Independent and D+I+R ratings are marginal. We, therefore, conjecture that aggregate ratings from partisan identities do not lead to significant improvement over the Independent identity, which already exhibits relatively low political bias.
Discussion
In this paper, we systematically audit eight widely used LLMs from three major providers to test their ability to discern credible information sources from low-credibility ones. For most news sources tested in our experiments, the LLMs could provide ratings as instructed. However, larger models tend to refuse to rate less popular sources due to insufficient knowledge, while smaller models tend to hallucinate responses. Comparing the LLM ratings, these models exhibit a high level of agreement, despite being trained by different providers. Conversely, the models only moderately correlate with human experts.
It remains unclear how the LLMs acquire this capacity. Based on the explanations provided by the models alongside the ratings, we hypothesize that the models summarize descriptions of the given news sources from their training data and generate ratings accordingly. For instance, when encountering high-credibility sources, the LLMs often note that these are well-established and reputable news websites. This could explain the high correlation among the LLMs, as they likely have common training data (Liu et al. 2024).
Focusing on news sources with clear political leanings in the US context, we find that assigning partisan identities to LLMs constantly steers their ratings to favor sources with congruent leanings. This result indicates that LLMs can reflect the viewpoints of humans with different political ideologies (Argyle et al. 2023; Simmons 2023). In their default configurations, LLMs exhibit a liberal bias: they are more likely to assign higher-than-expected credibility ratings for left-leaning sources compared to right-leaning ones. LLMs identified as Independents are the least biased, but still show a weak liberal bias. These findings align with previous studies indicating that many LLMs have left-leaning tendencies (Santurkar et al. 2023).
Our auditing has some limitations. There are different approaches to prompt the LLMs, which might yield different outcomes. For instance, one could employ a binary classification approach or ask LLMs to rank two sources at a time (Qin et al. 2024). We were unable to test all of these approaches. Additionally, our study focuses on only eight representative models, a tiny portion of LLMs on the market. Given the rapid development in the field, new models with different behaviors will emerge soon.
Despite these limitations, we believe our findings provide valuable insights into the current state-of-the-art LLMs and could be generalized to other models considering that most results are consistent regardless of the model size and provider. Given that these LLMs have been deployed in widely used systems, the findings have important implications.
Our experiments show that LLMs have the potential to serve as an affordable and accessible reference for source credibility ratings. Such ratings are vital for researchers to study the dissemination and prevalence of misinformation online (Lazer et al. 2018; Lin et al. 2023). With the ability to provide contextual information and actively answer user questions, LLMs might also be adopted by the general public as scalable media literacy tools to investigate the credibility of news sources and perform fact-checking (Pennycook and Rand 2019; Hoes, Altay, and Bermeo 2023). On the one hand, the accessibility and low cost of LLMs make them an appealing alternative to human rating services. On the other hand, our results indicate that the accuracy of LLMs is not perfect, calling for further analysis and comparison with other misinformation intervention methods (Kozyreva et al. 2024). Additionally, LLMs may be vulnerable to manipulation. Bad actors could pollute LLM training data by creating fake news source review websites.
Our findings also highlight the challenges of using LLMs as information curators. Although not confirmed by the providers, leaked system prompts of popular LLMs suggest that they are configured to retrieve information according to the quality and credibility of sources when searching the Internet, along with other criteria such as the relevance to user questions and the freshness of the information.101010github.com/jujumilk3/leaked-system-prompts However, it is unclear how the LLMs understand and implement these criteria in production.
A famous failure story came from Google in 2024, when its AI overview feature produced bizarre answers such as recommending that people eat rocks.111111bbc.com/news/articles/cd11gzejgz4o According to Google, these odd responses originated from low-quality information sources,121212blog.google/products/search/ai-overviews-update-may-2024 yet their AI models appeared to overlook it. Although these errors may have a limited negative impact, they highlight the risks of using LLMs for information curation in high-stakes contexts such as public health and elections.
Our findings provide valuable insights about such risks. We show that LLMs may lack information about many information sources, especially unpopular ones. However, it is inevitable for LLMs to encounter unfamiliar sources when curating information for users, especially when they search for misinformation topics (Boyd and Golebiewski 2018; Aslett et al. 2024). Under these circumstances, forcing the models to produce summaries or responses could inadvertently amplify lesser-known low-credibility sources. Even when the models know the information sources, including the highly popular ones, their assessment of credibility still deviates from human judgment. This inaccuracy might suppress credible sources and amplify low-credibility ones. When dealing with political information, the inherent bias of the LLMs in the default configuration might lead to distorted outcomes. The finding that assigning partisan identities to LLMs could induce corresponding biases suggests that LLMs tailored to user preferences and views might exacerbate echo chambers and polarization.
We also explore approaches to reduce political biases and increase the accuracy of the LLM ratings. Given the negative correlation between them, explicitly assigning an Independent identity to an LLM or mixing the ratings from different identities could help achieve both objectives simultaneously. However, even the unbiased ratings fail to align well with human judgment. Future models might consider explicitly integrating source credibility information into the training phases (Ouyang et al. 2022). When incorporating LLMs into larger systems such as search engines, one may also consider augmenting the models with external knowledge of the sources and instructing the models to give greater attention to credible ones.
Our study only scratches the surface when it comes to the impact of LLMs as information curators. Future studies should explore how LLMs handle information from different sources in more realistic scenarios, how other information selection criteria, such as relevance and freshness (Sun et al. 2023), affect the outcomes, and how to mitigate emerging manipulation techniques (Pfrommer et al. 2024). At the same time, more studies on how humans interact with AI-augmented systems (Sharma, Liao, and Xiao 2024; DeVerna et al. 2024) and how information processed by AI influences humans are needed (Jakesch et al. 2023; Simchon, Edwards, and Lewandowsky 2024).
Appendix A Appendix
Complete Prompts
In the default configuration, the following prompts are used in our experiments.
System prompt:
System prompt You are an assistant to determine the credibility of websites.
Task instruction:
Task instruction Rate the website’s credibility: {domain}, on a scale between 0 and 1, where 0 means very low credibility and 1 means very high credibility. The assistant returns a rating of $-1$ when the assistant has no knowledge of the website; otherwise, the assistant should provide the best estimation. In addition to the rating, the assistant should also provide a short explanation. The output should be in JSON format and follow the following schema. Only return the JSON object, nothing else. {‘url’: ‘example.com’, ‘rating’: 0.5, ‘explanation’: ‘The example website is known to post credible content.’}
Note that we specifically instruct the models to return the results in JSON format for easy parsing. We use the JSON mode for GPT and Gemini models to ensure their LLM responses are in valid JSON format. Unfortunately, the Llama models hosted by together.ai do not support JSON mode. They are usually able to follow the instructions and return the responses in JSON format, but extra cleaning and re-formatting are sometimes required to correctly parse the responses.
When testing the impact of political identities, we modify the system prompt:
Political identity assignment You are an assistant to determine the credibility of websites. You identify as {identity} on the US political spectrum.
In the experiments, {identity} is replaced with the one of “a Democrat,” “an Independent,” and “a Republican.” The task instruction remains the same.
Response Examples
In this section, we provide some response examples from the LLMs. Table 1 lists the LLM ratings of reuters.com. Table 2 lists the LLM ratings of aldf.com.
Table 1: Ratings of reuters.com by different LLMs.
Model Rating Explanation
GPT-4o mini 1 Reuters is a well-established international news organization known for its commitment to factual reporting and journalistic integrity. It has a strong reputation for accuracy and is widely regarded as a reliable source of news.
GPT-4o 1 Reuters is a well-established international news organization known for its accurate and unbiased reporting. It is widely respected in the journalism community and has a long history of providing reliable news.
GPT-4 0.95 Reuters is a well-established and highly reputable global news organization known for its commitment to accurate and unbiased reporting.
Llama 3.1 8B 0.9 Reuters is a well-established and reputable news agency with a long history of providing accurate and unbiased reporting.
Llama 3.1 70B 0.95 Reuters is a well-established and reputable news agency with a long history of providing accurate and unbiased reporting. It is widely considered a trusted source of news and information.
Llama 3.1 405B 0.9 Reuters is a well-established and reputable international news agency with a long history of providing accurate and unbiased reporting.
Gemini 1.5 Flash 0.9 Reuters is a well-established and reputable news agency known for its journalistic integrity and fact-checking practices.
Gemini 1.5 Pro 1 Reuters is a highly reputable international news agency known for its journalistic integrity and factual reporting.
Table 2: Ratings of aldf.com by different LLMs. Five models believe the website belongs to the Animal Legal Defense Fund (aldf.org).
Model Rating Explanation
GPT-4o mini 0.8 The Animal Legal Defense Fund (ALDF) is a well-established organization focused on animal rights and legal advocacy. It is recognized for its credible efforts in promoting animal welfare and has a strong reputation in the field.
GPT-4o -1 The assistant has no knowledge of the website aldf.com.
GPT-4 -1 I do not have enough information to assess the credibility of aldf.com.
Llama 3.1 8B 0.8 The Animal Legal Defense Fund (ALDF) is a well-established and reputable organization that provides accurate and reliable information on animal law and advocacy.
Llama 3.1 70B 0.8 The Animal Legal Defense Fund (ALDF) is a well-established and reputable organization that advocates for animal rights and welfare. Their website provides credible information on animal law and related topics.
Llama 3.1 405B 0.8 The Animal Legal Defense Fund (aldf.org) is a well-established organization with a clear mission and a history of advocating for animal welfare through the legal system. Their website provides credible information on animal law and related topics.
Gemini 1.5 Flash -1 I have no information about this website.
Gemini 1.5 Pro 1 ALDF.org is a highly reputable website of the Animal Legal Defense Fund, a well-known non-profit organization.